OpenStudy (mathgirl321):

Solve the proportion. 7 x - = - 45 9 a. 1.4 c. 3.8 b. 2.6 d. 5.5

OpenStudy (anonymous):

AAAAAAAAAAAAAAAAAAAA

OpenStudy (anonymous):

Cross multiply and solve. :) \[\frac{ 7 }{ 45 } = \frac{ x }{ 9 }\] Then multiply: 9*7 = 45 *x Divide: x = \[\frac{ 63 }{ 45 }\] So, x = A, 1.4 Hope that helped!